OpenStudy (anonymous):

6/14-x = 8/14+x solve the equation for x

OpenStudy (anonymous):

\[\frac{ 6 }{ 14 }- x = \frac{ 8 }{ 14 } + x\] First arrange like terms.

OpenStudy (anonymous):

i bet it is \[\frac{6}{14-x}=\frac{8}{14+x}\] but i could be wrong

OpenStudy (anonymous):

@satellite73 \[\frac{ 6-14x }{ 14 } = \frac{ 8 +14x}{ 14 }\] But its better if you arrange the like terms i.e. \[\frac{ 6 }{ 14 } - \frac{ 8 }{ 14 } = x + x\]